40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 25 15:43:29 2016, ericq, Update, SUS, OSEM issues SRM_OSEMS.pngPRM_LR.png
    Reply  Thu Aug 25 20:07:35 2016, gautam, Update, SUS, OSEM issues - maybe resolved? 
    Reply  Tue Sep 6 18:21:13 2016, rana, Update, SUS, OSEM issues PRMswing.png
       Reply  Tue Sep 6 20:52:42 2016, gautam, Update, SUS, OSEM issues P9060254.JPG
          Reply  Wed Sep 7 17:23:26 2016, gautam, Update, SUS, PRM LR fixed for now IMG_3208.JPGIMG_3211.JPG
             Reply  Tue Oct 11 00:43:50 2016, ericq, Update, SUS, PRM LR problematic again PRM_LR_Problem.pngsince_fix.pnglast10.png
                Reply  Tue Oct 11 10:15:04 2016, Steve, Update, SUS, PRM LR problematic again PRMalfoiled.png
                   Reply  Tue Oct 11 13:30:49 2016, gautam, Update, SUS, PRM LR problematic again PRM_LR_decay.png
                      Reply  Wed Oct 12 13:34:28 2016, gautam, Update, SUS, PRM LR problematic again 08.png
Message ID: 12472     Entry time: Tue Sep 6 18:21:13 2016     In reply to: 12441     Reply to this: 12475
Author: rana 
Type: Update 
Category: SUS 
Subject: OSEM issues 

I looked at the PRM free swing spectra. The modes look like they're at the right frequencies, so pointing more and more towards a LED or satellite box issue.

  7/2011 9/2016
POS 0.993 0.94
PIT 0.612 0.60
YAW 0.833 0.76
SIDE 0.999 0.993

Some of the frequencies have changed between the 2011 in-vac measurement and our 2016 in-air measurement, but that seems within usual parameters.

Attachment 1: PRMswing.png  122 kB  | Hide | Hide all
PRMswing.png
ELOG V3.1.3-